# CMoE: Fast Carving of Mixture-of-Experts for Efficient LLM Inference

### v1

@@ IV-A Shared and Routed Experts Grouping
Cluster Assignment Let $T_{i,p}^{t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p=1}^{N_{r}}T_{i,p}\cdot d_{i,p}$ (11)
s.t. $\sum_{i=1}^{N_{r}\cdot m}T_{i,p}=m,p=1,\dots N_{r}$
$\sum_{p=1}^{N_{r}}T_{i,p},i=1,\dots,N_{r}\cdot m$
$T_{i,p}\geq 0,i=1,\dots,N_{r}\cdot m,p=1,\dots N_{r}.$
Cluster Update Update $\hat{\mathbf{c}}_{p}^{t+1}$ as follows:
@@
The steps are terminated until $\hat{\mathbf{c}}_{p}^{t+1}=\hat{\mathbf{c}}_{p}^{t}$, $\forall p=1,\dots,N_{r}$.
We include the distance value $d_{i,j}$ into the cost function so that the balanced $K$-means clustering makes the intra-expert distances low and inter-expert distances high. However, the cluster assignment we defined above is an unbalanced assignment problem ($N_{r}\cdot m>m$) and cannot directly be solved by existing algorithms for balanced assignment. Therefore, we reduce it to a balanced assignment by extending the distance matrix $\mathbf{D}$ as follows:
@@
Balanced Assignment Let $T_{i,p^{\prime}}^{{}^{\prime},t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T^{\prime}}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}\cdot d_{i,p^{\prime}}^{ext}$ (13)
s.t. $\sum_{i=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,p^{\prime}=1,\dots N_{r}\cdot m$
$\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,i=1,\dots,N_{r}\cdot m$
$T^{\prime}_{i,p^{\prime}}\geq 0,i=1,\dots,N_{r}\cdot m,p^{\prime}=1,\dots N_{r}\cdot m,$
where $d_{i,p^{\prime}}^{ext}=d_{i,p}$ if $p\cdot m<p^{\prime}\leq(p+1)\cdot m$.
And the balanced cluster $\hat{\mathbf{c}}_{p}^{t+1}$ can be updated as follows:
@@
where $k_{0}=m(p-1)+1$. Drawing on the Jonker-Volgenant algorithm [32], this problem can be addressed as a reduced assignment problem in each step of the $K$-means algorithm, with a complexity of $O(n^{3})$.
The final solution $T^{\prime}$ gives us the optimized strategy to group the routed experts. We get the selection indices set $S_{N_{r},p}$, $p=1,\dots,N_{r}$, for each routed expert $E^{r}_{p}$:
@@
We then build the weights $\mathbf{W}_{up}^{r,p}\in\mathbb{R}^{d\times m}$, $\mathbf{W}_{gate}^{r,p}\in\mathbb{R}^{d\times m}$ and $\mathbf{W}_{down}^{r,p}\in\mathbb{R}^{m\times d}$ of each routed expert $E_{p}^{r}$ as in Equation 9. And the output of $E_{p}^{r}$ is given by:
@@
Suppose the MoE activates $N_{k}$ experts out of the $N_{r}$ routed experts. We expect the router to choose the routed experts with top $N_{k}$ scores. Therefore, we modify the MoE-version FFN as in (IV-A),
@@
$g_{i}$ =$=\left\{\begin{array}[]{rl}1,&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k}),\\ 0,&\text{otherwise},\end{array}\right.$ (17)
s $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=G(\textbf{x}),$ (18)
@@
IV-B Training-free Router Construction
We now present the training-free router network $G$ for CMoE. Unlike previous works, which either built the router from scratch or intuitively used hidden features as initialization, we formulate the router construction as a minimization problem and then develop an algorithm with analysis to construct the router by approximating the optimal solution.
Given the same input embedding $\mathbf{x}$, the output of original output of the dense FFN, i.e. $F(\mathbf{x})$ in (1), is equivalent to the sum of the output of all the experts in $F_{MoE}$:
$F(\mathbf{x})$ $=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}E_{i}^{r}(\textbf{x}).$ (19)
The only difference between (19) and $F_{MoE}(\mathbf{x})$ in (IV-A) is the expert score $\mathbf{g}$, which is obtained from the TopK selection of the output of $G$. Therefore, to preserve important knowledge captured by the original dense FNN, $G$ can be constructed to enforce $F_{MoE}(\mathbf{x})$ to be close to $F(\mathbf{x})$ by solving the minimization problem (IV-B),
$\arg\min_{G}|F_{MoE}(x;G)-F(x)|$
=\!= $\!\arg\min_{G}|\sum^{N_{r}}_{i=1}(g_{i}-1)E_{i}^{r}(\textbf{x})|=\!\!\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|,$
where $S_{de}\!=\!\{i:s_{i}\!\notin\!\text{TopK}(\{s_{i}|1\!\leq\!j\!\leq\!N_{r}\},N_{k})\}$ and $|S_{de}|\!=\!N_{r}\!-\!N_{k}$, and the problem becomes constructing the $G$ to minimize the absolute sum of the output of deactivated routed experts. Note that we have made a hypothesis in (4) that the output/sparsity of $F(x)$ is highly related to the norm/sparsity of $\mathbf{h}$, which is the same for the expert outputs. Based on (3) and (4), we reformulate (IV-B) as in (20):
$\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|$ (20)
@@ IV-B Training-free Router Construction
$s_{\sigma(1)}\leq s_{\sigma(2)}\leq\dots\leq s_{\sigma(N_{r})}\mbox{ and}$
$\bar{\mathbf{h}}^{r}_{\sigma(1)}\leq\bar{\mathbf{h}}^{r}_{\sigma(2)}\leq\dots\leq\bar{\mathbf{h}}^{r}_{\sigma(N_{r})},$ (21)
@@
$S_{de}$ =$=\{\sigma(1),\dots,\sigma(N_{r}-N_{k})\}$
=$=\{i:s_{i}\notin\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k})\}.$ (23)
Finally, we pay attention to the hidden state $\mathbf{h}^{r}$ of any routed expert, which is the output by the neurons that we grouped in Section IV-A, where the cluster is centered to the centroid $\hat{\mathbf{c}}$. We denote the neuron in each cluster that has the lowest $L_{2}$ distance to the centroid as the representative neuron:
@@
$G(\mathbf{x})$ $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=\left[h^{r}_{R_{1}},h^{r}_{R_{2}},\cdots,h^{r}_{R_{N_{r}}}\right]$
$\approx\left[\bar{\mathbf{h}}^{r}_{1},\mathbf{h}^{r}_{2},\cdots,\mathbf{h}^{r}_{N_{r}}\right],$ (26)
which is hence an approximate solution for the original problem introduced in (IV-B).
IV-C Differentiable Routing and Load-balancing
Though we have constructed a well-designed router in Section IV-B, it is not differentiable since each expert score $g_{i}$ is a constant, as shown in (IV-A), hindering further alignment and performance recovery. Therefore, we introduce a learnable parameter $\mathbf{u}$ when computing the expert scores as follows,
$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (28)
$\mathbf{s}^{\prime}$ $=\text{Softmax}(\mathbf{s}),\ \textbf{u}=[u_{1},u_{2},\dots,u_{N_{r}}]$ (29)
@@ IV-C Differentiable Routing and Load-balancing
For MoE models, load-balancing is crucial to guarantee computational efficiency, especially expert parallelism in LLMs serving. As in DeepSeek-V3 [4], we use the auxiliary-loss-free load balancing by introducing a bias term $\mathbf{b}$ before the TopK selection:
$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}+b_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (32)
$\mathbf{b}$ $=[b_{1},b_{2},\dots,b_{N_{r}}].$ (33)
@@
V Experiments
TABLE I: Comparison of perplexity results. Randomly select samples from WikiText-2 training dataset for calibration and training.
Method Type LLaMA-2-7B LLaMA-3-8B Training-free Fine-tuning Training-free Fine-tuning WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 Dense - 5.27 7.27 - - 6.14 9.44 - - LLaMA-MoE A2E8 nan nan 468.00 2660.68 nan nan 988.20 7521.83 LLaMA-MoE A4E16 nan nan 540.62 2690.68 nan nan 1094.24 7758.91 CMoE S1A1E8 60.86 135.61 12.76 32.12 162.74 324.71 21.16 64.03 CMoE S1A3E16 89.19 180.65 13.84 33.56 262.85 465.09 22.97 72.34 CMoE S2A2E16 62.30 136.12 12.73 32.37 143.38 284.19 21.01 65.57
CMoE is implemented based on Hugging Face Transformers [33] together with Pytorch [34]. The experiments are conducted on one NVIDIA H800 PCIe 80GB graphics card with CUDA Driver 12.6. We randomly select samples from WikiText-2 training dataset [35] as calibration and fine-tuning data. We use only 8 examples with 2,048 sequence length as calibration for calculating the hidden state $\mathbf{H}$ in (6). We set $K_{a}=10$ for the activation status record, which sounds counter-intuitive but works best in practice. For lightweight fine-tuning, we run $1$ epoch using the Adam optimizer [36] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95. We also employ LoRA [37] with a rank of 8 and $lora\_alpha=32$. We fine-tune all the models including the baselines with $2,048$ samples. We set different initial learning rates for the score scale $\mathbf{u}$ and other parameters, i.e. $0.001$ and $5.95e^{-5}$, respectively. We set the bias update speed $\gamma$ to $0.001$.
V-A Main Results
We compare CMoE with the up-to-date baseline LLaMA-MoE [11], in which the neurons are randomly split, and continual pre-training is carried out with additional router networks. The experiments cover both training-free and lightweight fine-tuning versions. We design these experiments to demonstrate the remarkable post-training performance of CMoE. we demonstrate the results on two different datasets, i.e. WikiText-2 [35] and C4 [38]. The baseline models are chosen as LLaMA-2-7B and LLaMa-3-8B.
Language Modeling Performance. We evaluate the perplexity of the constructed Mixture-of-Experts (MoE) models in Table I. We use abbreviations to denote the composition of experts. For instance, ‘S2A2E16’ represents 2 shared experts, 2 activated routed experts, and a total of 16 experts and the total activation ratio is $(2+2)/16=25\%$. Our findings indicate that in the training-free scenario, the ppl of our baseline LLaMA-MoE on both LLaMA-2-7B and LLaMA-3-8B are NaN across different datasets. In contrast, the perplexity of the proposed CMoE can be effectively controlled. Moreover, after fine-tuning, we observe that the perplexity can be reduced to as low as 12.73 when the activation ratio is 25%, which corresponds to a sparsity of 75%. Additionally, the CMoE model outperforms all state-of-the-art (SOTA) structured pruning methods [9, 22, 23].
Downstream Tasks Performance. We also evaluate the performance of the constructed MoE models on various downstream tasks. We evaluate on the following benchmarks: 32-shot BoolQ [39], 0-shot PIQA [40], 0-shot SciQ [41], 5-shot Winogrande [42], 25-shot ARC-Challenge [43], and 10-shot HellaSwag [44]. The results are presented in table II, where we denote fine-tuning or not with ‘FT’. In all activation ratio configurations, we find that the proposed CMoE outperforms the LLaMA-MoE in accuracy across a diverse set of downstream tasks, on both training-free and fine-tuning scenarios.
To illustrate the performance of the models, we select the SciQ dataset for zero-shot testing and the BoolQ dataset for few-shot testing as representative examples. When the activation ratio is set at a fixed value of 25% (corresponding to a sparsity rate of 75%), after fine-tuning, the LLaMA-MoE model can only achieve an accuracy equivalent to 21.2% of that of the relevant dense model on the SciQ dataset and 45.31% on the BoolQ dataset. Given that the perplexity (ppl) of the LLaMA-MoE model is Not a Number (NaN) in the training-free scenario, we do not report its accuracy in this case. In contrast, even without fine-tuning (i.e., in the training-free scenario), the proposed CMoE model demonstrates superior performance, attaining an accuracy equivalent to 56.3% of that of the dense model on the SciQ dataset and 54% on the BoolQ dataset. After fine-tuning, the performance of the CMoE model is further enhanced, with the accuracy reaching 76.59% on the SciQ dataset and 74.3% on the BoolQ dataset (relative to dense model).
V-B Ablation Studies
We conduct ablation studies with LLaMA-2-7B and data randomly selected from WikiText-2 training datasets.
Figure 4: Trade-off between Model Performance and Construction Time with Increasing Training Data.

### v2

@@ 3.1 Shared and Routed Experts Grouping
Cluster Assignment. Let $T_{i,p}^{t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T}$ $\sum_{i=1}^{mN_{r}}\sum_{p=1}^{N_{r}}T_{i,p}\cdot d_{i,p}$ (7)
s.t. $\sum_{i=1}^{mN_{r}}T_{i,p}=m,\forall p\in\{1,\dots,N_{r}\};\;\sum_{p=1}^{N_{r}}T_{i,p}=1,\;\forall i\in\{1,\dots,mN_{r}\};\;T_{i,p}\geq 0,\forall p,i.$
Cluster Update. Update $\hat{\mathbf{c}}_{p}^{t+1}$ as follows:
@@
The stopping criterion is $\hat{\mathbf{c}}_{p}^{t+1}=\hat{\mathbf{c}}_{p}^{t}$, $\forall p=1,\dots,N_{r}$. We include the distance value $d_{i,j}$ into the cost function so that the balanced $K$-means clustering makes the intra-expert distances low and inter-expert distances high. However, the cluster assignment defined above is an unbalanced assignment problem ($mN_{r}>m$) and cannot be solved directly by existing algorithms for balanced assignment. Therefore, we reduce it to a balanced assignment by extending the distance matrix $\mathbf{D}$ as follows:
@@
Balanced Assignment. Let $T_{i,p^{\prime}}^{{}^{\prime},t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T^{\prime}}$ $\sum_{i=1}^{mN_{r}}\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}\cdot d_{i,p^{\prime}}^{ext}$ (9)
s.t. $\sum_{i=1}^{mN_{r}}T^{\prime}_{i,p^{\prime}}=1,\forall p^{\prime}\in\{1,\dots,mN_{r}\};\;\sum_{p^{\prime}=1}^{mN_{r}}T^{\prime}_{i,p^{\prime}}=1,\forall i\in\{1,\dots,mN_{r}\};\;T^{\prime}_{i,p^{\prime}}\geq 0,\forall I,p^{\prime},$
where $d_{i,p^{\prime}}^{ext}=d_{i,p}$ if $mp<p^{\prime}\leq m(p+1)$. The balanced cluster $\hat{\mathbf{c}}_{p}^{t+1}$ can be updated as follows:
@@
where $k_{0}=m(p-1)+1$. The final solution $T^{\prime}$ gives us the optimized strategy to group the routed experts. We get the selection indices set $S_{N_{r},p}$, $p=1,\dots,N_{r}$, for each routed expert $E^{r}_{p}$:
@@
We then build the weights $\mathbf{W}_{up}^{r,p}\in\mathbb{R}^{d\times m}$, $\mathbf{W}_{gate}^{r,p}\in\mathbb{R}^{d\times m}$ and $\mathbf{W}_{down}^{r,p}\in\mathbb{R}^{m\times d}$ of each routed expert $E_{p}^{r}$ as in (LABEL:eq:expert_w). And the output of $E_{p}^{r}$ is given by:
@@
Suppose the MoE activates $N_{k}$ experts out of the $N_{r}$ routed experts. We expect the router to choose the routed experts with top $N_{k}$ scores. Therefore, we modify the MoE-version FFN as in (3.1),
@@
$g_{i}=\left\{\begin{array}[]{rl}1,&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k}),\\ 0,&\text{otherwise},\end{array}\right.\mbox{ and }\;\textbf{s}=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=G(\textbf{x}),$ (13)
@@
3.2 Training-free Router Construction
Unlike previous works, which either built the router from scratch or intuitively used hidden features as initialization, we formulate the router construction as a minimization problem and then develop an algorithm with analysis to construct the router by approximating the optimal solution.
Given the same input embedding $\mathbf{x}$, the output of original output of the dense FFN is equivalent to the sum of the output of all the experts in $F_{MoE}$: $F(\mathbf{x})=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}E_{i}^{r}(\textbf{x})$. The only difference roots in the expert score $\mathbf{g}$, which is obtained from the TopK selection of the output of $G$. Therefore, to preserve important knowledge captured by the original dense FNN, $G$ can be constructed to enforce $F_{MoE}(\mathbf{x})$ to be close to $F(\mathbf{x})$ by solving the minimization problem (14),
$\arg\min_{G}|F_{MoE}(x;G)-F(x)|=\arg\min_{G}|\sum^{N_{r}}_{i=1}(g_{i}-1)E_{i}^{r}(\textbf{x})|=\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|,$ (14)
where $S_{de}\!=\!\{i:s_{i}\!\notin\!\text{TopK}(\{s_{i}|1\!\leq\!j\!\leq\!N_{r}\},N_{k})\}$ and $|S_{de}|\!=\!N_{r}\!-\!N_{k}$, and the problem becomes constructing the $G$ to minimize the absolute sum of the output of deactivated routed experts.
Note that we have made a hypothesis in (3) that the output/sparsity of $F(x)$ is highly related to the norm/sparsity of $\mathbf{h}$, which is the same for the expert outputs. We reformulate (14) as in (15):
@@ 3.2 Training-free Router Construction
$s_{\sigma(1)}\leq s_{\sigma(2)}\leq\dots\leq s_{\sigma(N_{r})}\mbox{ and}\ \ \ \bar{\mathbf{h}}^{r}_{\sigma(1)}\leq\bar{\mathbf{h}}^{r}_{\sigma(2)}\leq\dots\leq\bar{\mathbf{h}}^{r}_{\sigma(N_{r})},$ (16)
@@
$S_{de}=\{\sigma(1),\dots,\sigma(N_{r}-N_{k})\}=\{i:s_{i}\notin\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k})\}.$ (18)
The neuron closest to the centroid in each cluster is considered as the representative neuron:
@@
$G(\mathbf{x})=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=\left[h^{r}_{R_{1}},h^{r}_{R_{2}},\cdots,h^{r}_{R_{N_{r}}}\right]\approx\left[\bar{\mathbf{h}}^{r}_{1},\mathbf{h}^{r}_{2},\cdots,\mathbf{h}^{r}_{N_{r}}\right],$ (21)
which is hence an approximate solution for the original problem introduced in (14).
3.3 Differentiable Routing and Load-balancing
Though we have constructed a well-designed router in Section 3.2, it is not differentiable since each expert score $g_{i}$ is a constant, as shown in (3.1), hindering further alignment and performance recovery. Therefore, we introduce a learnable parameter $\mathbf{u}$ when computing the expert scores as follows,
$g_{i}\!\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.\mathbf{s}^{\prime}=\text{Softmax}(\mathbf{s}),\;\&\;\textbf{u}=[u_{1},u_{2},\dots,u_{N_{r}}],$ (23)
@@ 3.3 Differentiable Routing and Load-balancing
For load balancing, we introduce a bias term $\mathbf{b}$ before the TopK selection as follows:
$g_{i}\!\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}+b_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.\mbox{ and }\;\mathbf{b}=[b_{1},b_{2},\dots,b_{N_{r}}].$ (26)
@@
4 Experiments
CMoE is implemented using Hugging Face Transformers wolf2019huggingface and PyTorch paszke2019pytorch . We use samples from WikiText-2 merity2016pointer for calibration and initial fine-tuning, and SlimPajama cerebras2023slimpajama for extended fine-tuning experiments.
Calibration: We randomly select 8 examples (2048 sequence length) from WikiText-2 to compute activation statistics for neuron grouping and initial router construction (Eq. 27). We set $K_{a}=10$ for the activation status record, which sounds counter-intuitive but works best in practice.
Lightweight Fine-tuning (2k): We fine-tune using LoRA hu2021lora (rank 8, alpha 32) on 2,048 WikiText-2 samples for 1 epoch. The optimizer is Adam kingma2014adam ($\beta_{1}=0.9,\beta_{2}=0.95$). We use different learning rates for the router scale parameter (0.001) and other LoRA parameters (5.95e-5). The load balancing bias update speed $\gamma=0.001$.
Extended Fine-tuning (SlimPajama): To assess performance recovery potential, we conduct further fine-tuning on subsets of the diverse SlimPajama cerebras2023slimpajama dataset (up to 1.2B tokens), using similar LoRA settings and run 1 epoch.
4.1 Main Results
We compare CMoE against several baselines on PPL, downstream task performance and inference three aspects. (1) Dense Models: The original Llama-2 7B and Llama-3 8B checkpoints. (2) LLaMA-MoE llama-moe : Random neuron split followed by continual pre-training (we evaluate its training-free state and after our lightweight 2k fine-tune for comparison). (3) LLaMA-MoE-v2 llama-moe-v2 : Initialize router networks based on hidden features input and do importance-based neuron partition by measuring the loss change of neuron pruned.
Our main CMoE configurations vary the activation ratio (Total Activated Experts / Total Experts) while keeping the ratio of shared-to-routed experts 1:1. For example, $25\%$ activation uses S1A1E8 (1 shared + 1 active routed / 8 total). For baselines, we also use 8 total experts in default.
Table 1: Language modeling perplexity (PPL) on WikiText-2 and C4 validation sets. Lower is better. CMoE variants vary activation ratio (fixed 1:1 shared/routed). ‘TF’: Training-Free, ‘FT’: LoRA fine-tuned on 2k WikiText-2 samples.
Method Act. Llama-2 7B Llama-3 8B WikiText-2 C4 WikiText-2 C4 TF FT TF FT TF FT TF FT Dense 100% 5.27 - 7.27 - 6.14 - 9.44 - LLaMA-MoE 25% nan 468.00 nan 2660.68 nan 988.20 nan 7521.83 LLaMA-MoE-v2 25% >10k 17.72 >10k 55.10 >60k 2754.80 >50k 5300.97 CMoE 25% 62.30 12.73 136.12 32.37 143.38 21.01 284.19 65.57 LLaMA-MoE 75% >60k 322.20 >50k 2061.88 >500k 970.29 >400k 7079.15 LLaMA-MoE-v2 75% >10k 8.68 7358.15 19.76 >10k 1411.23 >10k 2843.39 CMoE 75% 7.02 5.69 10.98 10.21 10.16 8.26 17.35 17.06